Identify the center and radius of x²+y²-8x+6y+16=0.Then sketch the graph.


1
Expert's answer
2021-11-23T12:59:32-0500

"x^2+y^2-8x+6y+16=0"

"(x^2-8x+16)+(y^2+6y+9)-9=0"

"(x-4)^2+(y+3)^2=3^2"


"O(4, -3)", "R=3"